OpenStudy (anonymous):

Find all real roots. 1.fourth roots of 625 2.fifth roots of -243 3.fourth roots of 256y^8

OpenStudy (anonymous):

@wolf1728

OpenStudy (solomonzelman):

\[625=5^4\]\[-243=(-6)^3\]\[256y^8=2^8y^8=(2y)^8\]GO....

OpenStudy (solomonzelman):

(those are tips, not answers)

OpenStudy (anonymous):

lol okay thankyou hold on le tme try and finish it

OpenStudy (solomonzelman):

sure.

OpenStudy (anonymous):

for the first one would I raise both sides to the fourth power to cancel out the 5^4?

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

YES.

OpenStudy (anonymous):

I got a really large number..? I don't think that's right..?

OpenStudy (solomonzelman):

\[\sqrt[4]{625}=\sqrt[4]{5^4}=5\]

OpenStudy (solomonzelman):

Wasn't this what you meant?

OpenStudy (anonymous):

oh yeah... sorry calculation error

OpenStudy (solomonzelman):

\[\sqrt[5]{-243}=\sqrt[5]{(-3)^5}=?\]

OpenStudy (solomonzelman):

\[\sqrt[4]{256y^8}=\sqrt[4]{4^4y^8}=4~\sqrt[4]{y^8}=4~\sqrt[4]{(y^2)~^4}=4y^2\]

OpenStudy (solomonzelman):

Makes sense?

OpenStudy (anonymous):

wait, sorry was the first one for #2 and the second one for #3?

OpenStudy (anonymous):

but yeah it makes sense though

OpenStudy (wolf1728):

If you are familiar with logarithms, here is the first problem: log(625) = 2.7958800173 Dividing that by 4 we get, .69897 Get 10 and raise it to the .69897 power (or looking up the anti-log) 10^.69897 equals 5

OpenStudy (anonymous):

I have no idea what logarithms are?

OpenStudy (wolf1728):

Okay, well that is how you would do that if you knew them. I suppose if you have a math calculator you could solve it. Does it have a key that reads y^x

OpenStudy (anonymous):

it would, if I had my calculator with me at the moment...:l

OpenStudy (wolf1728):

Do you have Excel or OpenOffice on your computer?

OpenStudy (anonymous):

yes I do

OpenStudy (wolf1728):

Very good. Well switch to that then in any cell enter =625^.25

OpenStudy (anonymous):

entered

OpenStudy (wolf1728):

There should be an answer right in the same box

OpenStudy (wolf1728):

Press "Enter" or press the down arrow

OpenStudy (anonymous):

I didn't get anything.. weirdly enough..

OpenStudy (wolf1728):

When you move out of the cell does anything remain in the cell?

OpenStudy (wolf1728):

If you know how to copy and paste copy this equation =625^.25 paste it into any cell, then click Enter or press an arrow key

OpenStudy (anonymous):

that's what I did and nothing happened.. but usually it does, I've used excel before.. idk what's going on.

OpenStudy (wolf1728):

It should work. I'm using Open Office and it works just fine. Excel works exactly the same.

OpenStudy (anonymous):

nvm i think I got it

OpenStudy (wolf1728):

A five should appear in that cell right?

OpenStudy (anonymous):

Yeah.

OpenStudy (wolf1728):

Okay - now you can use Excel instead of your calculator!

OpenStudy (anonymous):

Yes I can. Thankyou! :)

OpenStudy (wolf1728):

We can move to the fifth root of -243 now.

OpenStudy (anonymous):

ok

OpenStudy (wolf1728):

Even though Excel can help you, sometimes it might fail when you are looking for negative roots. When the exponent is even you get imaginary numbers but when it is odd, it is possible to get roots. So, let's find the 5th root of 243 - that's right POSITIVE 243

OpenStudy (wolf1728):

If you are waiting for the formula it is =243^.2 We are raising 243 to the 2/10 power which equals 1/5 - the fifth root.

OpenStudy (anonymous):

no sorry I was trying to find the original problem. I found out it now ,but thanks

OpenStudy (wolf1728):

Anyway, try the fifth root of 243

OpenStudy (anonymous):

243^.2

OpenStudy (anonymous):

oops sorry wrong one I got 3

OpenStudy (wolf1728):

That is it - 3 is correct!

OpenStudy (anonymous):

yaay! thanks

OpenStudy (wolf1728):

But remember we have to find the fifth root of -243, so the answer is actually -3. To be sure multiply -3 5 times. =-3*-3*-3*-3*-3

OpenStudy (anonymous):

-243

OpenStudy (wolf1728):

Pretty good huh?

OpenStudy (wolf1728):

That leaves us with 4th root of 256y^8

OpenStudy (anonymous):

mhm

OpenStudy (wolf1728):

We really don't need Excel for this. 256 = 16 * 16 and 16= 4 * 4 So the fourth root of 246 is 4 the 4th root of y^8 is ? (Do you know?)

OpenStudy (wolf1728):

Made a mistake in ABOVE problem - should be 256 for both numbers.

OpenStudy (anonymous):

y^3?

OpenStudy (anonymous):

y^3 =y^8?

OpenStudy (wolf1728):

no because if you multiply y^3 four time you'd get y^12

OpenStudy (anonymous):

sorry I meant y^2

OpenStudy (wolf1728):

Yes y² is the answer. Now for the entire answer the 4th root of 256y^8 is ?

OpenStudy (anonymous):

4y^2

OpenStudy (wolf1728):

Yes !!! Very good

OpenStudy (anonymous):

ThankYou!!! :)

OpenStudy (anonymous):

If I as you a question and show you my work for it, Can you tell me what I did wrong please? If you're not busy.

OpenStudy (anonymous):

*ask

OpenStudy (wolf1728):

Well okay sure.

OpenStudy (anonymous):

for...\[\frac{ 6x }{ x+4 }=\frac{ 7x+4 }{ x+4 }\] I Multipilied both sides by (x+4) to cancel out the denominators and got 6x=7x+4. Then I subtracted 7 on both sides and got -1x=4. Finally, I divied both sided by -1 and got x=-4.

OpenStudy (wolf1728):

What I would do is cross-multiply and get (x+4) / (x+4) = 7x + 4 / x+4 7x + 4 / x+4 = 1

OpenStudy (anonymous):

but then wouldn't you multiply both sides by x+4 to cancel out the denominator?

OpenStudy (wolf1728):

I don't know. I guess that would give you 6x = 7x +4 Whereas I got 7x + 4 / 6x = 1 and I forgot to type in the 6 a little further back. So far, I'd say we have the same equation.

OpenStudy (anonymous):

did you get the same answer as me though?

OpenStudy (wolf1728):

Looking at your answer, for 6x = 7x +4 You can't subtract 7 from both sides Is that where you get the -1 ? 6x -7 = -1?

OpenStudy (anonymous):

hm we got the same answer...

OpenStudy (wolf1728):

But 6x -7 does not equal -1

OpenStudy (wolf1728):

I think my German Shepherd wants to go out for a walk - I might be gone for a while (15 minutes maybe)

OpenStudy (anonymous):

oh yeah sorry, my brain's not thinking straight! & Lucky! You have a German Shepherd! alright i'll try to this on my own 'til them. :)

OpenStudy (anonymous):

In case, I got to sleep by the time you comeback.. can you tell me what I did wrong in these.. I'll look over them tomorrow morning. @wolf1728

OpenStudy (anonymous):

If $2000 is invested at 4% interest compounded monthly the value of the investment after t years is given by 2000(12.04/12) ^12t. What is the value of the investment after 3.5 years?...I actually didn't know how to do this problem at all.

OpenStudy (anonymous):

\[\frac{ n^2-6n-27 }{ n^2-10n+9 }=\frac{ (n-9)(n+3) }{ (n-9)(n-1) }=\frac{ (n+3) }{ (n-1)}\] this is what I got, but I don't know what I did wrong same with the next one that's similar.

OpenStudy (anonymous):

\[\frac{ 3n^3-36n^2+105n }{ 49n-7 }=\frac{ 3n(n^2-12n+35) }{ 7(7n-1) }=\frac{ 3n(n-7)(n-5) }{ 7(7n-1) }\]

OpenStudy (wolf1728):

I see that you are now offline, so I guess I'll work on this alone. Looking again at this: 6x = 7x +4 Multiply both sides by 1/6 x = (7/6)x + 4/6 -(1/6) x = 4/6 -x = 4 OR x = -4 Which works for 6x = 7x + 4 But when in the original equation 6x / (x+4) = 7x+4 / (x+4) -4 produces division by zero.

OpenStudy (wolf1728):

Okay, I'm really good at compound interest. $2,000.00 invested at 4% compounded monthly, value after 't' years is? After 3.5 years? The first thing to do is convert 4% compounded monthly to annual rate. annual rate = ((1 + (.04/12))^12) -1 annual rate = .0407415429 or 4.07415429% Present Value = principal * (1.0407415429)^years Present Value = 2,000.00 * (1.0407415429)^years For 3.5 years = 2,000.00 * (1.0407415429)^3.5 After 3.5 years = 2,000.00 * 1.1500060278 After 3.5 years = 2,300.01 Here's a good link for compound interest: http://www.1728.org/compint.htm

OpenStudy (wolf1728):

n²−6n−27 n²−10n+9 (n−9)(n+3)(n−9)(n−1)=(n+3)(n−1) Well I looked at this one and the other one but I didn't find anything wrong with either. I hope the compound interest answer helps you out.